OpenStudy (anonymous):

Short Answer. Add the following polynomials: (4x^2 – 3x + 2) + (5x^3 – 7x^2 – 9)

OpenStudy (anonymous):

5x^3-3x^2 -3x -7

OpenStudy (anonymous):

(2x + 5y) + (3x – 2y) = 2x + 5y + 3x – 2y = 2x + 3x + 5y – 2y = 5x + 3y

OpenStudy (anonymous):

that is wrong

OpenStudy (anonymous):

y

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

look at the question

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

you are wrong